Find the maximum or minimum value of the function.
f(x)= -x^2/3+2x+6 (it's -x^2 over 3 then +2x+6)

I thought that you have to use -b/2a and then f(x)=-b/2a

Answer by stanbon(75887)

Find the maximum or minimum value of the function.
f(x)= -x^2/3+2x+6 (it's -x^2 over 3 then +2x+6)
----
a = (-1/3) ; b = 2 ; c = 6
----
I thought that you have to use -b/2a and then f(x)=-b/2a
Max or min occurs at x = -b/(2a) = -2/(2*(-1/3)) = -2/(-2/3) = 3
----
f(3) = (-1/3)(3^2) + 2(3) + 6
= -3+6+6
---
= 9
===========
Cheers,
Stan H.
===========